Suppose you enter an elevator at a certain floor. The elevator moves up 6 floors, down 4 floors, and up 2 floors. You are then at floor 7. At which floor did you initially enter the elevator?

Answer by ankor@dixie-net.com(22740)

Suppose you enter an elevator at a certain floor.
The elevator moves up 6 floors, down 4 floors, and up 2 floors.
You are then at floor 7. At which floor did you initially enter the elevator?
:
Let x = floor that you initially entered
:
x + 6 - 4 + 2 = 7
x + 4 = 7
x = 7 - 4
x = 3rd floor